Write an ARM assembly language subroutine (named n
Answer #1
nfibo proc
  PUSH EBP          ; save previous frame pointer
  MOV  EBP, ESP     ; set current frame pointer

  MOV  EAX, [EBP+8] ; get argument N
  CMP  EAX, 1       ; N<=1?
  JA   Recurse      ; no, compute it recursively
  MOV  ECX, 1       ; yes, Fib(1)--> 1
  JMP  exit

Recurse:
  DEC  EAX          ; = N-1
  MOV  EDX, EAX     ; = N-1
  PUSH EDX          ; save N-1
  PUSH EAX          ; set argument = N-1
  CALL nfibo    ; compute Fib(N-1) to ECX
  POP  EAX          ; pop N-1
  DEC  EAX          ; = N-2
  PUSH ECX          ; save Fib(N-1)
  PUSH EAX          ; set argument = N-2
  CALL nfibo    ; compute Fib(N-2) to ECX
  POP  EAX          ; = Fib(N-1)
  ADD  ECX, EAX     ; = Fib(N-1)+FIB(N-2)
exit:
  MOV  ESP,EBP      ; reset stack to value at function entry 
  POP  EBP          ; restore caller's frame pointer
  RET               ; and return